OpenStudy (anonymous):

I need help A steel lift column in a service station is a solid cylinder 4.0 m long and 0.20 m in diameter. Young's modulus for this steel is 20 × 1010 N/m2. By what distance does the column compress when a 5000-kg truck is on it?

OpenStudy (anonymous):

Young's modulus, E, can be calculated by dividing the tensile stress by the extensional strain in the elastic (initial, linear) portion of the stress–strain curve E is the Young's modulus (modulus of elasticity) F is the force exerted on an object under tension; A0 is the original cross-sectional area through which the force is applied; ΔL is the amount by which the length of the object changes; L0 is the original length of the object. E = (F/A0)/(ΔL/L0) since you look for ΔL ΔL/L0 = (F/A0)/E ΔL= (L0*(F/A0))/E hope this helps

OpenStudy (anonymous):

also F = m*g m is mass g represents gravitational acceleration constant wich is ussually taken as either 10 m/s^2 or 9.8 m/s^2

OpenStudy (anonymous):

Would I need to use g in solving the problem?

OpenStudy (anonymous):

I see got it 5000 * 9.8

OpenStudy (anonymous):

i think so, since you need force wich is not given in question part and only way to solve it is to take that force applied to surface is simply directed towards centre of earth. I mean, this is how i imagine it